Solve the differential equation with the given initial condition. y + 2xy = 8x y(0) = 0 y(x) =
Answer #1

da Given, y+ 2xy = 87, 46)=0 iedy + (2x) Y = 8X — which is linear differen -tial equation of the form dy + py=a Herre P=28,given, Y0=0 froros * 460) = 44C. (1) tie=1) 0 = 4+0 [c=-4] Hence, 4(x) = 4+ (-4ed (0) - [y(x) = 4 (1-2) which is required sol